# appendixtocaptAppendix

### v3

@@ 3.2.1 Environment and Experiment Setup
To accommodate environmental complexity, we modify our random walk dynamics to incorporate spatial constraints imposed by obstacles and boundaries. For any location $x$, transitions to neighboring locations $y$ that are prohibited by the obstacles setting $p(y|x,t=1)=0$ for these forbidden transitions. Consequently, the self-transition probability is adjusted to maintain normalization $p(x|x,t=1)=1-|N(x)|\cdot p_{\text{move}}$. The key difference between open field and obstacle-containing environments lies in the value of $N(x)$. In open fields, $N(x)$ consistently includes all 8 neighboring locations (except at boundaries). With obstacles, $N(x)$ represents only the subset of accessible neighboring locations from position $x$, which could be smaller than $8$. We maintain $p_{move}=1/9$ across all environments for consistency. Despite these modifications to the underlying transition dynamics, our learning approach remains consistent. We compute the normalized transition probabilities $q(y|x,t)$ based on these constraint-respecting dynamics and train our model to learn the embedding function $h(x,t)$ using the same objective function as in the open field.
Figure 2: Place Cells in Complex Maze Environments. (A) Path planning through obstacle-containing environments. (B) Topologically-informed transition kernels $q(y|x,t)$ with gradient fields. (C) Spatial coverage despite environmental complexity. (D) Remapping with environmental modification.
@@ 3.2.2 Navigation Efficiency
h Environment Success SPL ($\uparrow$) Open field $100\%$ $0.991\pm 0.04$ U-shape $100\%$ $0.919\pm 0.25$ S-shape $100\%$ $1.392\pm 0.98$ Four-room $100\%$ $1.519\pm 1.17$
@@
To quantify efficiency, we use Success weighted by inverse Path Length (SPL) [2], where values near $1.0$ indicate near-optimal paths and values exceeding $1.0$ suggest our method found shortcuts the Bug algorithm missed. Details of SPL calculation appear in Appendix G. In simpler environments (U-shape), our method performs comparably to the oracle-enhanced Bug algorithm. However, in complex environments (S-shape and Four-room), our approach significantly outperforms the baseline, discovering more efficient paths that the Bug algorithm fails to identify even with oracle assistance.
@@ 3.2.3 Transition and Learned Profile in Complex Environments
To evaluate the adaptive capabilities of our approach, we examine the transition kernel $q(y|x,t)$ and the resulting place cell activation profiles $h(x,t)$ in environments containing complex obstacles. Figure 2 (B) presents the normalized transition probability distributions $q(y|x,t)$ for strategically selected locations, with a large scale to reveal long-range spatial relationships. The diffraction-like patterns, where probability flow encounters obstacles, reveal how spatial information propagates through available pathways while respecting environmental constraints.
@@
We visualized the collective place field distributions across all experimental environments in Figure 2 (C). Despite increasing environmental complexity, the learned place cell population maintains comprehensive spatial coverage throughout accessible regions. The overlapping receptive fields are distributed uniformly across the space. This consistent coverage property demonstrates the model’s robust ability to develop effective spatial representations regardless of environmental geometry, a critical feature for reliable navigation in diverse settings.
@@ 3.2.4 Remapping Properties
4 Discussion on Theta-Phase and Grid Cells
There are two major omissions in our current experiments on place cells so far:
(1) Theta-phase procession based on angle-phase duality. The hippocampus not only encodes spatial relationships but also organizes temporal dynamics through theta phase precession, where place cells fire at specific phases of the theta rhythm as an animal traverses their fields [47, 55]. We may extend our population embedding framework to model this phenomenon. First, the learned place fields tile the environment. Let $\mu_{i}=\arg\max_{x}h_{i}(x,t)$, i.e., the center of the place field for cell $i$. The theta phase of cell $i$ can be defined in terms of the angle between the current $h(x,t)$ and $h(\mu_{i},t)$. As $x$ approaches $\mu_{i}$, the angle changes from $\pi/2$ to 0, and as $x$ moves away from $\mu_{i}$, the angle changes from 0 to $\pi/2$. Appendix E provides details.
(2) Grid cells as 2D coordinates and linear basis. For scientific reductionism, we focus on place cells without incorporating grid cells. A natural extension would explore the relationship $h(x,t)=W(t)g(x)$, where $g(x)$ represents the vector of grid cell activations and $W(t)$ is a learned transformation matrix, potentially unifying our framework with the complementary roles of place and grid cells in spatial representation [41]. In fact, the brain has no access to $x$ explicitly, and $g$ plays the role of $x$. Thus for proper path planning, such as calculating the gradient of $q(y|x,t)$, $g$ must preserve distance and angle in $x$, that is, $g$ must be conformal to $x$ [67, 66, 20, 21]. Moreover, $g$ should encode nonlinear function of $x$, such as $h(x,t)$, via linear projection $W(t)g$ [20]. Essentially we treat $h(x,t)$ as a neural network with $g$ serving as input embedding of $x$. Appendix F provides the formulation, and we will conduct experiments in future work.
In addition, we have not integrated sensory inputs for spatial navigation[38, 46], nor have we studied navigation in a 3D or dynamic environment [31, 68]. Moreover, we have not studied the on-line learning of place cells over sequential experiences to model temporal memories [14, 49].
5 Related Work
5.1 Place Cell Models and Representations
@@
5.2 Multi-scale Spatial Representations
@@
5.3 Navigation and Path Planning
@@
5.4 Inner Product Spaces in Neural Representation
@@
5.5 Heat Equation and Diffusion Models in Neuroscience
@@
6 Conclusion
This paper reconceptualizes hippocampal place cells, shifting from individual place fields to proximity-preserving neural embeddings of the whole population, akin to viewing a tree rather than leaves. These embeddings encode multi-scale spatial proximity defined by multi-time random walk transitions, where the inner product $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$ approximates the normalized transition probability, with $q(x|x,t)=1$ and $\|h(x,t)\|=1$. The embeddings $h(x,t)$ arise from spectral analysis of $q(y|x,t)$ and non-negative matrix factorization, leading to probability-geometry duality. Non-negative constraint ensures localized place fields that collectively tile the environment. The time parameter $\sqrt{t}$ defines the spatial scale, mirroring hippocampal multi-scale organization. Despite being purely random, the random walk’s transitions encode geodesic distance for short $t$ and topological connectivity for long $t$, capturing environmental structure.
Our framework employs gradient ascent on $q(y|x,t)$ with adaptive scale selection, minimizing the Euclidean distance $\|h(x,t)-h(y,t)\|^{2}$ for trap-free, smooth trajectories. We term this straight-forward path planning, as it follows the gradient of Euclidean distance in the embedding space, simplifying complex physical navigation. Efficient matrix squaring ($P_{2t}=P_{t}^{2}$) builds global transitions from local ones, enabling preplay-like shortcut prediction. Extensible to theta-phase precession through phase-angle duality, our model unifies spatial navigation, multi-scale coding, and temporal dynamics, offering a biologically plausible, computationally efficient framework for complex environments.
Our framework is about learning for planning, instead of learning for reaction as in reinforcement learning. Our method only requires data from local exploration, with matrix squaring performing trajectory stitching and systematic generalization. Compared to the dynamics of the so-called world model, our embeddings encode a far more informative cognitive map of the state space to support straight-forward planning. In our future work, we shall explore manipulation tasks using our method.
[…]
Appendix A Spectral Analysis of Multi-Time Random Walk
This appendix provides a detailed spectral analysis of the multi-time random walk transition kernel, establishing connections between our random walk model and position embeddings.
A.1 Eigendecomposition of the Transition Matrix (Background)
@@
$P_{1}=Q\Lambda Q^{T}$ (16)
with orthogonal $Q$ ($Q^{T}Q=I$) and diagonal $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$, where $0\leq\lambda_{i}\leq 1$, and we assume the random walk is irreducible and aperiodic. The multi-step transition matrix is:
$P_{1}^{t}=Q\Lambda^{t}Q^{T}$ (17)
@@
A.2 Position Embeddings from Spectral Decomposition
@@
A.3 Normalization and Non-Negative Embeddings
@@
$h_{\text{spec}}(x,t)=\frac{H(x,t)}{\|H(x,t)\|}=\frac{H(x,t)}{\sqrt{p(x|x,t)}}$ (21)
@@
$\langle h_{\text{spec}}(x,t),h_{\text{spec}}(y,t)\rangle=\frac{p(y|x,t)}{\sqrt{p(x|x,t)\cdot p(y|y,t)}}=q(y|x,t)$ (22)
However, $h_{\text{spec}}(x,t)$ may contain negative components, which conflicts with the biological constraint that neural firing rates must be non-negative. This is where Horn’s theorem becomes relevant.
A.4 Non-Negative Matrix Factorization (Background)
@@
For our normalized transition matrix $Q_{t}=[q(y|x,t)]$, which is symmetric with non-negative entries, Horn’s theorem guarantees the existence of a non-negative matrix $H_{t}$ such that $Q_{t}=H_{t}H_{t}^{T}$. The rows of $H_{t}$ provide our desired non-negative position embeddings $h(x,t)$.
Appendix B Heat Equation with Reflecting Boundaries (Background)
@@
B.1 From Discrete Random Walk to Reflecting Heat Equation
We begin with a discrete random walk on a two-dimensional integer grid. For simplicity, we assume the one-step transition $p(y|x,t=1)$ of the random walk is to move to one of 4 nearest neighbors with $p_{\rm move}=1/4$. We assume this simplest $p(y|x,t=1)$ in our theoretical derivations in all the relevant sections in the Appendix. Similar results can be obtained for more general $p(y|x,t=1)$, with a different diffusion coefficient $\alpha$.
@@ B.1 From Discrete Random Walk to Reflecting Heat Equation
$x=i\cdot dx,\;\;y=j\cdot dx,\;\;t=k\cdot dt$ (23)
@@
B.2 Derivation of the Heat Equation
@@
$\frac{\partial p(x,y,t)}{\partial t}=\alpha\nabla^{2}p(x,y,t)$ (30)
B.3 Reflecting Boundary Conditions
Our discrete random walk enforces reflecting boundary conditions, preventing probability flow into obstacles, mirroring hippocampal obstacle avoidance [38]. Consider a boundary point $(i,j)$ with an obstacle at $(i+1,j)$, so the number of valid neighbors is $N(i,j)=3$ (points $(i-1,j)$, $(i,j+1)$, $(i,j-1)$). The self-transition probability is $p_{\text{stay}}=1-\frac{N(i,j)}{4}=\frac{1}{4}$, redistributing probability that would flow to the obstacle back to $(i,j)$.
@@ B.3 Reflecting Boundary Conditions
B.4 Connection to Geodesic Distance
Varadhan’s formula [60] establishes a deep relationship between the heat kernel and geodesic distance. For a complete Riemannian manifold $M$ with heat kernel $p(x,y,t)$, Varadhan proved that:
$\lim_{t\to 0}-4t\log p(x,y,t)=d_{g}^{2}(x,y)$ (34)

### v4

@@ 3.2.1 Environment and Experiment Setup
To accommodate environmental complexity, we modify our random walk dynamics to incorporate spatial constraints imposed by obstacles and boundaries. For any location $x$, transitions to neighboring locations $y$ that are prohibited by the obstacles setting $p(y|x,\tau=1)=0$ for these forbidden transitions. Consequently, the self-transition probability is adjusted to maintain normalization $p(x|x,\tau=1)=1-|N(x)|\cdot p_{\text{move}}$. The key difference between open field and obstacle-containing environments lies in the value of $N(x)$. In open fields, $N(x)$ consistently includes all 8 neighboring locations (except at boundaries). With obstacles, $N(x)$ represents only the subset of accessible neighboring locations from position $x$, which could be smaller than $8$. We maintain $p_{\rm move}=1/9$ across all environments for consistency. Despite these modifications to the underlying transition dynamics, our learning approach remains consistent. We compute the normalized transition probabilities $q(y|x,\tau)$ based on these constraint-respecting dynamics and train our model to learn the embedding function $h(x,\tau)$ using the same objective function as in the open field.
Figure 2: Place Cells in Complex Maze Environments. (A) Path planning through obstacle-containing environments. (B) Topologically-informed transition kernels $q(y|x,\tau)$ with gradient fields. Target $y$ is marked as a red point. (C) Randomly sampled place cell profiles at multiple spatial scales. (D) Remapping with environmental modification.
@@ 3.2.2 Navigation Efficiency
Environment Success SPL ($\uparrow$)
Open field $100\%$ $0.991\pm 0.04$
U-shape $100\%$ $0.919\pm 0.25$
S-shape $100\%$ $1.392\pm 0.98$
Four-room $100\%$ $1.519\pm 1.17$
To quantify efficiency, we use Success weighted by inverse Path Length (SPL) [28], where values near $1.0$ indicate near-optimal paths and values exceeding $1.0$ suggest our method found shortcuts the Bug algorithm missed. Details of SPL calculation appear in Appendix J. In simpler environments (U-shape), our method performs comparably to the oracle-enhanced Bug algorithm. However, in complex environments (S-shape and Four-room), our approach significantly outperforms the baseline, discovering more efficient paths that the Bug algorithm fails to identify even with oracle assistance.
Table 2: Path planning in Four-Room environment.
Metrics Bug (w/ Oracle) Bug (w/o Oracle) A* Search Random Walk Our Method
Success (%) $100\%$ $14\%$ $100\%$ $2\%$ $100\%$
SPL ($\uparrow$) $0.66\pm 0.51$ $0.15\pm 0.00$ $1.08\pm 0.50$ $\pm 0.00 1.001.00$
We also compared our method with several different baselines in the Four-Room environment, with each trial limited to 50,000 steps. The results in Table 2 show that baseline approaches like the Bug algorithm (without oracle guidance) and Random Walk fail to solve most of the challenging scenarios.
@@ 3.2.3 Transition and Learned Profile in Complex Environments
To evaluate the adaptive capabilities of our approach, we examine the transition kernel $q(y|x,\tau)$ and the resulting place cell activation profiles $h(x,\tau)$ in environments containing complex obstacles. Figure 2 (B) presents the normalized transition probability distributions $q(y|x,\tau)$ for strategically selected locations, with a large scale to reveal long-range spatial relationships. The diffraction-like patterns, where probability flow encounters obstacles, reveal how spatial information propagates through available pathways while respecting environmental constraints.
@@
We visualized place cell profiles at multiple spatial scales across complex maze environments in Figure 2 (C). Each panel displays a small subset of non-overlapping place cells at different scales $\tau$, emphasizing the emergent sparsity and localized nature of individual receptive fields. Despite increased environmental complexity, the learned place cell population maintains comprehensive spatial coverage throughout accessible regions. At smaller scales, place fields are tightly localized; at larger scales, fields broaden while maintaining their localized tiling structure. This multi-scale coverage property demonstrates the model’s robust ability to develop effective spatial representations regardless of environmental geometry, a critical feature for reliable navigation in diverse settings.
@@ 3.2.4 Remapping Properties
4 Conclusion
This paper reconceptualizes hippocampal place cells as population embeddings approximating multi-step random walk transition kernels, where $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$, offering a biologically plausible model of spatial navigation [29, 30]. The time parameter $\sqrt{\tau}$ defines a multi-scale representation, mirroring dorsoventral place field scaling [5], adaptability [4], and remapping [31]. Gradient ascent on $q(y|x,\tau)$ with adaptive scale selection produces trap-free trajectories, guided by boundary avoidance, diffraction-like passage navigation, aligning with hippocampal cognitive maps [11, 9, 3]. Efficient matrix squaring ($P_{2\tau}=P_{\tau}^{2}$) computes global transitions from local ones ($P_{1}$) without past trajectory memorization, enabling preplay-like shortcut detection [10, 7, 8]. Bridging connectionist models [12, 13] and cognitive map theories [2, 9], our framework captures hippocampal navigation’s dynamic properties, providing a scalable, computationally efficient model for complex environments.
Acknowledgments
Y. W. is partially supported by NSF DMS-2415226, DARPA W912CG25CA007 and research gift funds from Amazon and Qualcomm.
Appendix A Related Work
A.1 Place Cell Models and Representations
@@
A.2 Multi-scale Spatial Representations
@@
A.3 Navigation and Path Planning
@@
A.4 Successor Representation
Our work is closely related to the Successor Representation (SR) [49, 50], as both frameworks build upon powers of a transition matrix to construct a representation of space. The SR is typically defined as a discounted sum, $M=\sum_{k=0}^{\infty}(\gamma T)^{k}$, which yields the expected discounted future occupancy of all states from any given starting state, and the discount factor $\gamma$ implicitly sets a single, predictive temporal horizon. In contrast, our approach utilizes the discrete powers of the transition matrix directly, $P_{\tau}=P^{\tau}$, to explicitly model the transition probabilities at multiple, distinct time horizons, where the parameter $\tau$ corresponds directly to the time scale of a random walk. This multi-scale representation aligns more closely with the observed functional gradient of place field sizes along the hippocampal dorsoventral axis.
The primary novelty of our framework lies in learning a biologically-constrained, non-negative matrix factorization of these transition kernels, where vector embeddings $h(x,\tau)$ are learned such that their inner product reconstructs the transition probability, $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$. A crucial theoretical insight is that the combination of this inner product objective with non-negativity and orthogonality constraints for distant locations gives rise to emergent sparsity; the embeddings are forced to have disjoint support sets, providing a geometric explanation for the localized firing fields of place cells without explicit regularization. This multi-scale architecture confers significant functional advantages, enabling an adaptive gradient-ascent-based navigation policy that selects the optimal scale $\tau^{*}$ at each step to generate smooth, trap-free trajectories. This approach is not only more scalable, but also offers a more direct navigation mechanism that bypasses the need for explicit value function computation common in SR-based reinforcement learning agents.
A.5 Inner Product Spaces in Neural Representation
@@
A.6 Heat Equation and Diffusion Models in Neuroscience
@@
Appendix B Limitations
B.1 Lack of Sensory Integration
Our current model does not incorporate sensory inputs for spatial navigation[11, 2]. Real hippocampal place cells integrate multimodal sensory information including visual landmarks, self-motion cues from path integration, and boundary detection. This sensory integration is fundamental to how place cells form their spatial receptive fields and anchor representations to environmental features.
Without sensory processing, our model cannot explain how place fields emerge during initial exposure to novel environments. In our framework, position embeddings $h(x,\tau)$ are learned from pre-defined transition probabilities that already encode environmental structure, rather than being built up through sensory experience. We also cannot model perceptual anchoring, where place cells maintain stable firing relative to visual landmarks, or how sensory conflicts (such as navigation in darkness) affect place cell stability. Future extensions incorporating vision-based inputs, self-motion integration, and boundary vector cell signals would significantly enhance biological plausibility.
B.2 Limited to 2D Static Environments
Our experiments and analyses are restricted to two-dimensional static environments, whereas biological spatial navigation operates in three-dimensional space and continuously changing contexts[5, 59]. Studies in flying bats have revealed volumetric place cells encoding 3D positions, and terrestrial animals navigating multi-level structures require three-dimensional representations. Extension to 3D introduces complications our framework does not address: vertical movements and gravity-dependent asymmetries, different place field scaling across dimensions, and substantially increased computational complexity. Dynamic environments would require time-varying transition probabilities and continuous updating mechanisms our framework does not provide.
B.3 Online Sequential Learning
Another limitation is our lack of online learning over sequential experiences to model temporal dynamics and memory consolidation[7, 8]. Our batch optimization learns $h(x,\tau)$ across all spatial locations simultaneously, fundamentally differing from how biological place cells develop through experience. Biological place cell formation is incremental: initial weak or unstable spatial tuning gradually sharpens through repeated exploration via activity-dependent synaptic plasticity. Our framework does not capture this gradual emergence from unstructured initial conditions. Our matrix squaring $P_{2\tau}=P_{\tau}^{2}$ achieves global integration instantaneously based on environmental structure, not through iterative learning over sleep-wake cycles analogous to biological replay.
Despite these limitations, our framework offers valuable theoretical insights into computational principles underlying hippocampal spatial navigation, particularly the role of multi-scale transition probabilities, emergence of sparsity from geometric constraints, and the connection between random walk dynamics and cognitive maps.
Appendix C Spectral Analysis of Multi-Step Random Walk
This appendix provides a detailed spectral analysis of the multi-step random walk transition kernel, establishing connections between our random walk model and position embeddings.
C.1 Eigendecomposition of the Transition Matrix (Background)
@@
$P_{1}=Q\Lambda Q^{T},$ (13)
with orthogonal $Q$ ($Q^{T}Q=I$) and diagonal $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$, where $0\leq\lambda_{i}\leq 1$, where we assume the random walk is irreducible and aperiodic. The multi-step transition matrix is:
$P_{1}^{\tau}=Q\Lambda^{\tau}Q^{T}$ (14)
@@
C.2 Position Embeddings from Spectral Decomposition
@@
C.3 Normalization and Non-Negative Embeddings
@@
$h_{\text{spec}}(x,\tau)=\frac{H(x,\tau)}{\|H(x,\tau)\|}=\frac{H(x,\tau)}{\sqrt{p(x|x,\tau)}}$ (18)
@@
$\langle h_{\text{spec}}(x,\tau),h_{\text{spec}}(y,\tau)\rangle=\frac{p(y|x,\tau)}{\sqrt{p(x|x,\tau)\cdot p(y|y,\tau)}}=q(y|x,\tau)$ (19)
However, $h_{\text{spec}}(x,\tau)$ may contain negative components, which conflicts with the biological constraint that neural firing rates must be non-negative. This is where Horn’s theorem becomes relevant.
C.4 Non-Negative Matrix Factorization (Background)
@@
For our normalized transition matrix $Q_{\tau}=[q(y|x,\tau)]$, which is symmetric with non-negative entries, Horn’s theorem guarantees the existence of a non-negative matrix $H_{\tau}$ such that $Q_{\tau}=H_{\tau}H_{\tau}^{T}$. The rows of $H_{\tau}$ provide our desired non-negative position embeddings $h(x,\tau)$.
Appendix D Heat Diffusion with Reflecting Boundaries (Background)
@@
D.1 From Discrete Random Walk to Reflecting Heat Equation
We begin with a discrete random walk on a two-dimensional integer grid. For simplicity, we assume the one-step transition $p(y|x,\tau=1)$ of the random walk is to move to one of 4 nearest neighbors with $p_{\rm move}=1/4$. We assume this simplest $p(y|x,\tau=1)$ in our theoretical derivations in all the relevant sections in the Appendix. Similar results can be obtained for more general $p(y|x,\tau=1)$, with a different diffusion coefficient $\alpha$.
@@ D.1 From Discrete Random Walk to Reflecting Heat Equation
$x=i\cdot dx,\;\;y=j\cdot dx,\;\;\tau=k\cdot dt$ (20)
@@
D.2 Derivation of the Heat Equation
@@
$\frac{\partial p(x,y,\tau)}{\partial\tau}=\alpha\nabla^{2}p(x,y,\tau)$ (27)
D.3 Reflecting Boundary Conditions
Our discrete random walk enforces reflecting boundary conditions, preventing probability flow into obstacles, mirroring hippocampal obstacle avoidance [11]. Consider a boundary point $(i,j)$ with an obstacle at $(i+1,j)$, so the number of valid neighbors is $N(i,j)=3$ (points (i1,j)(i-1,j), $(i,j+1)$, (i,j1)(i,j-1)). The self-transition probability is $p_{\text{stay}}=1-\frac{N(i,j)}{4}=\frac{1}{4}$, redistributing probability that would flow to the obstacle back to $(i,j)$.
@@ D.3 Reflecting Boundary Conditions
D.4 Connection to Geodesic Distance
Varadhan’s formula [16] establishes a deep relationship between the heat kernel and geodesic distance. For a complete Riemannian manifold $M$ with heat kernel $p(x,y,\tau)$, Varadhan proved that:
$\lim_{\tau\to 0}-4t\log p(x,y,\tau)=d_{g}^{2}(x,y)$ (31)